Thomas Toys Ltd. uses a periodic review inventory management system. One important item for the company is building blocks, which sell, on average, five sets per day. However, the standard deviation of demand is two sets per day. The company checks the status of inventory for building blocks every 8 days. When blocks are reordered from the supplier, it takes 12 days to be replenished. Thomas has just checked its inventory and found that it currently has 165 sets in stock. The company desires to maintain a 97.5 percent service level.

a. How many sets of building blocks should Thomas Toys order? (Negative answer should be indicated by a minus sign. Do not round intermediate calculations. Round your final answer to the nearest whole number.)
Answer #1

Maximum daily consumption= 5+2= 7 units

Maximum Lead time= 12 days of order receipt +8 days of stock check =20 days

Average daily consumption=5

Average lead time=12 days

Safety stock= (max. daily consumption X Maximum lead time)-(average daily consumption x average lead time) = (20*7)-(12*5)= 80 Units.

Service level = 97.5%

So safety stock becomes= 80*.975= 78

Current stock: 165

Therefore at present there is no need to order.